# On the first test of the Weak Equivalence Principle in low Earth orbit

### v4

@@ V Outliers, gaps and artificial data
Table 3: Synodic frequencies of the MICROSCOPE spacecraft relative to the Sun (left column; fewer digits are shown because it is approximated by the spin frequency) and the Earth (right column) up to the 5th harmonic. All frequencies are given in $\rm mHz$ and refer to spin mode V3 ($\nu_{{}_{spin}}=35\nu_{{}_{orb}}/2$) .
@@
An FFT analysis of large glitches (with $\rm SNR>3$) for a typical measurement session is reported in Fig. 7 of Ref. (30). It shows that the large glitches occur at all the frequencies listed in Table 3 at the level of $10^{-11}\,\rm ms^{-2}$ and above, reaching about $7\times 10^{-11}\,\rm ms^{-2}$ at the spin frequency, i.e. the synodic frequency relative to the Sun. Higher harmonics appear too, but they are less relevant to the experiment. In the logarithmic scale of the Figure each line appears as a double line because the two synodic frequencies are close to each other (see Table 3), except the line at the orbital frequency which is the difference of the two. However, in the inset of the same Figure showing a zoom in the region of the spin frequency, the scale is linear and the two synodic frequencies clearly appear as distinct lines.
In Ref.(30) it is concluded that glitches must be removed from the data because it is impossible to accurately model their effects which include one at the frequency of the signal. In the time series of each session glitches are identified as all outliers above $4.5\,\sigma$ from the moving average, some time is allowed after each outlier to account for its decay, and all data points identified in this way are removed, up to 35% in SUEP and up to 40% in SUREF. As a result, the time series of the remaining data is interspersed with gaps. Gaps are then filled with artificial reconstructed data so that the resulting time series can be analyzed in the frequency domain to compute the spectral density of each measurement session, hence the limit to WEP violation $\delta_{i}\pm\sigma_{i}$. By including all $19$ SUEP sessions the final limit $\eta=\delta_{fin}\pm\sigma_{fin}=(-1.5\pm 2.3)\times 10^{-15}$ is obtained (Ref.(4)), where:8
@@
and $\sigma_{fin}^{2}=1/\big{(}{\sum\limits_{i=1}^{i=19}\frac{1}{\sigma_{i}^{2}}}\big{)}$.
@@
The second issue is that the artificial data at the level of random platform noise which replace the outliers do contribute to increasing the integration time of each measurement session, hence to reducing the level $\delta_{i}$ of WEP violation of the session (Eq.6). Then, according to Eq.(11), the final result is also reduced. If this procedure were correct, the more outliers are present, the larger the fraction of gaps, the larger the fraction of artificial data, the better the WEP test of the session and the WEP test of the entire experiment. Which is clearly a paradox. No matter how artificial data are generated and/or manipulated, they should not introduce any physical information, i.e. their effect on the $\delta$ estimated in each session ought to be null.
An interesting comparison is reported in Ref.(30) between the FFT of the differential acceleration obtained from the original measured data, and the FFT after glitches were removed and the missing data were reconstructed. This was done for SUREF session #380 lasting 120 orbits. This session has been chosen because, with equal composition test masses, any effect at $\nu_{{}_{EP}}$ is spurious, and the low noise makes the effects of glitches well visible.
Fig. 10 of Ref.(30), shows the two plots, FFT of the measured data in black, of reconstructed data in red. All the double lines listed in Table 3 (plus the single line at $\nu_{{}_{orb}}$) are clearly visible (the tidal line at $2\nu_{{}_{EP}}$ has not been removed). It is apparent that the procedure used to eliminate the glitches only reduces their effects, including the one at $\nu_{{}_{EP}}$. In fact, in the final SUREF analysis this session, originally with $46\,\%$ of artificial data, has been split into two short sessions (of $46$ and $34$ orbits), each one with a small percentage of glitches ($7\,\%$ and $5\,\%$). Thus, most of the reconstructed data have been eliminated, which is a reasonable decision.
@@
Yet, in all cases $\delta_{i}$ has been obtained with the same procedure, eliminating the same outliers, generating the same gaps, filling the gaps with the same data, the difference being only the fake signal which is therefore obviously recovered. The problem is that $\delta_{i}$ itsef may be affected by the artificial data, and this fact is not addressed by the check.
@@
VI Conclusions
A space test of the weak equivalence principle in the field of the Earth has enormous potential for a leap forward in precision, by building on a strong driving signal and better instrument isolation from nearby masses in space than in any ground laboratory, hence easier control of systematic effects from local disturbances.29
@@ VI Conclusions
The experiment appears to be limited by thermal noise from internal damping in the grounding wires whose stiffness was predicted incorrectly in contrast with previous results from GOCE. Once in orbit, in order to reduce the higher level of noise, the spin rate of the satellite was raised to a considerably larger rate than planned, only to find an unexplained noise jump. We find jumps, up or down, also in sequential sessions at the same spin rate, which are inexplicable on the basis of the current knowledge of internal damping noise and the physical quantities it depends upon.
@@
Any future WEP experiment in space must prove by its very design that it will not suffer from glitches (GOCE showed the way), and must avoid introducing artificial data potentially compromising the analysis.
@@
Acknowledgements. Thanks are due to Giuseppe Catastini for many discussions on the heritage from ESA’s GOCE mission. While at Alcatel Alenia Space (now Thales Alenia Space), GC was the chief scientist responsible for the design, development and operation of the GOCE end-to-end software simulator, a crucial tool for a mission the performance of which could not be fully tested on the ground.

### v6

@@ 5 Outliers, gaps and artificial data
Table 3: Synodic frequencies of the Microscope spacecraft relative to the Sun, $\nu_{{syn\odot}}$ (left column; fewer digits are shown because it is approximated by the spin frequency) and the Earth, $\nu_{{}_{syn\oplus}}$ (right column) up to the 5th harmonic. All frequencies are given in $\rm mHz$ and refer to spin mode V3 ($\nu_{{}_{spin}}=N\nu_{{}_{orb}}/2$ with $N=35$).
@@
An FFT analysis of large glitches (with $\rm SNR>3$) for a typical measurement session is reported in Fig. 7 of  Bergé et al (2022a). It shows that the large glitches occur at all the frequencies listed in Table 3 at the level of $10^{-11}\,\rm ms^{-2}$ and above, reaching about $7\times 10^{-11}\,\rm ms^{-2}$ at the spin frequency, i.e. the synodic frequency relative to the Sun. Higher harmonics appear too, but they are less relevant to the experiment. In the logarithmic scale of the Figure each line appears as a double line because the two synodic frequencies are close to each other (see Table 3), except the line at the orbital frequency which is the difference of the two. However, in the inset of the same Figure showing a zoom in the region of the spin frequency, the scale is linear and the two synodic frequencies clearly appear as distinct lines.
In Bergé et al (2022a) it is concluded that glitches must be removed from the data because it is impossible to accurately model their effects which include one at the frequency of the signal. In the time series of each session glitches are identified as all outliers above $4.5\,\sigma$ from the moving average, some time is allowed after each outlier to account for its decay, and all data points identified in this way are removed, up to 35% in SUEP and up to 40% in SUREF. As a result, the time series of the remaining data is interspersed with gaps. Gaps are then filled with artificial reconstructed data so that the resulting time series can be analyzed in the frequency domain to compute the spectral density of each measurement session, hence the limit to WEP violation $\delta_{i}\pm\sigma_{i}$. By including all $19$ SUEP sessions the final limit $\eta=\delta_{fin}\pm\sigma_{fin}=(-1.5\pm 2.3)\times 10^{-15}$ reported in Touboul et al (2022a) is obtained, with:
@@
and $\sigma_{fin}^{2}=1/\big{(}{\sum\limits_{i=1}^{i=19}\frac{1}{\sigma_{i}^{2}}}\big{)}$ (see Touboul et al (2022b)).
@@
The second issue is that the artificial data at the level of random platform noise which replace the outliers do contribute to increasing the integration time of each measurement session, hence to reducing the level $\delta_{i}$ of WEP violation of the session (Eq.6), i.e. to improving the test. Then, according to Eq.(11), the final result of the whole experiment is also improved. If this procedure were correct, the more outliers are present, the larger the fraction of gaps, the larger the fraction of artificial data, the better the WEP test of the session and the WEP test of the entire experiment. Which is clearly a paradox. No matter how artificial data are generated and/or manipulated, they should not introduce any physical information, i.e. their effect on the $\delta$ estimated in each session ought to be null.
An interesting comparison is reported in Bergé et al (2022a) between the FFT of the differential acceleration obtained from the original measured data, and the FFT after glitches were removed and the missing data were reconstructed. This was done for SUREF session #380 lasting 120 orbits. This session has been chosen because, with equal composition test masses, any effect at $\nu_{{}_{EP}}$ is certainly spurious and moreover the low noise makes the effects of glitches well visible. Fig. 10 of Bergé et al (2022a) shows the two plots, FFT of the measured data in black, of reconstructed data in red. All the double lines listed in Table 3 (plus the single line at $\nu_{{}_{orb}}$) are clearly visible (the tidal line at $2\nu_{{}_{EP}}$ has not been removed). It is apparent that the effects of glitches, including the one at $\nu_{{}_{EP}}$, have been reduced but not eliminated. In fact, in the final SUREF analysis this session, originally with $46\,\%$ of artificial data, has been split into two short sessions (of $46$ and $34$ orbits), each one with a small percentage of glitches ($7\,\%$ and $5\,\%$). Thus, most of the reconstructed data have been eliminated, which is a reasonable decision.
@@
Yet, in all cases $\delta_{i}$ has been obtained with the same procedure, eliminating the same outliers, generating the same gaps, filling the gaps with the same data, the difference being only the fake signal which is therefore obviously recovered. The problem is that $\delta_{i}$ itself may be affected by the artificial data, and this fact is not addressed by the check.
@@
6 Conclusions
A space test of the weak equivalence principle in the field of the Earth has enormous potential for a leap forward in precision, by building on a strong driving signal and better instrument isolation from nearby masses in space than in any ground laboratory, hence easier control of systematic effects from local disturbances (Nobili and Anselmi (2018a)).
@@ 6 Conclusions
The experiment appears to be limited by thermal noise from internal damping in the grounding wires whose stiffness was predicted incorrectly (by two orders of magnitude) in contrast with previous results from GOCE. Once in orbit, in order to reduce the higher level of noise, the spin rate of the satellite was raised to a considerably larger rate than planned, and an unexplained noise jump was oserved. We find jumps, up or down, also in sequential sessions at the same spin rate, which are inexplicable on the basis of the current knowledge of internal damping noise and the physical quantities it depends upon.
@@
Any future WEP experiment in space must prove by its very design that it will not suffer from glitches (GOCE showed the way) and must avoid introducing artificial data potentially compromising the analysis.
@@
Acknowledgements.
Thanks are due to Giuseppe Catastini for many discussions on the heritage from ESA’s GOCE mission. While at Alcatel Alenia Space (now Thales Alenia Space), GC was the chief scientist responsible for the design, development and operation of the GOCE end-to-end software simulator, a crucial tool for a mission the performance of which could not be fully tested on the ground.
Conflict of Interest
The authors declare that they have no conflict of interest.